# LCDB 1.1: A Database Illustrating Learning Curves Are More Ill-Behaved Than Previously Thought

### v1

@@ 5.3 DA2: Can Feature Scaling Mitigate Ill-Behavior?
To understand the impact of feature scaling, we now compare the results across scaling techniques. Table 4 indicates that feature scaling marginally reduces the amount of ill-behavior. From Figure 4 we observe that for most learners, feature scaling does not resolve ill-behavior. The Sigmoid SVM becomes slightly more monotone and has fewer peaks, but is still significantly ill-behaved. While preprocessing does not reduce ill-behavior, the SVM absolute performance improves notably and training becomes more stable after feature scaling is applied (see Appendix E.2 for details). Nearest Centroid and some Naive Bayes models are the only models that become significantly more ill-behaved with feature scaling. The biggest reductions in ill-behavior occur for the Ridge classifier and MLP. Ridge almost becomes completely monotone and without peaks when using min-max scaling, but not with standard scaling. The MLP improves significantly when using standard scaling, largely resolving ill-behaviors, however, min-max scaling does not always help the MLP. We confirm LDA is insensitive to feature scaling, and we find the peak occurs when the training set size is approximately equal to the dimensionality (Appendix E.3) in line with peaking literature.
@@ 5.4 DA3: How Do Monotonicity and Convexity Violations Affect Curve Fitting?
[…]
@@ 6 Discussion and Conclusion
The shape analysis is challenging. Note that ill-behavior may change if curves are longer (especially dipping). It is also difficult to maintain statistical rigor and consistency; this is because the Bonferroni correction imposes different levels of stringency, for instance, when testing monotonicity (two anchors) and convexity (three anchors). Bonferroni is also quite pessimistic, and some subjective choices had to be made. For example, peaking can also be detected differently (see Appendix H), yet results are similar. An analysis using E-values [60] or controlling the false discovery rate [61] may alleviate inconsistency issues, but such an analysis is non-trivial, going beyond our main point.
@@
We have tried our best to make LCDB 1.1 fully reproducible, by using a docker container and fixed python package versions, yet we find that one LDA curve and several QDA curves are non-reproducible, likely due to numerical non-determinism of the singular value decomposition. For this reason, we have dropped the QDA learner from the analysis.
@@
The next step is to investigate whether ill-behavior persists under hyperparameter tuning, which we leave for future work, since collecting learning curve data with tuning is computationally expensive (LCDB 1.1 already required 600K CPU hours; see Appendix I). Moreover, hyperparameters in Scikit have reasonable defaults that are determined by the community, and as such the ill-behavior observed remain surprising and relevant, especially when persistent to different scalings.
@@
In conclusion, we introduce the Learning Curves Database 1.1 (LCDB 1.1). This database is more reproducible, of higher resolution, and has multiple types of preprocessing (with and without data-leakage), making it a valuable database for the community to study learning curves. Moreover, we carefully study ill-behavior and find that a significant amount, 14%, of the curves exhibit ill-behavior while some learners misbehave more frequently than others. Feature scaling rarely solves this problem and in some cases can make it worse. Lastly, we demonstrate the impact of ill-behavior on downstream tasks, underscoring the practical implications. We hope that LCDB 1.1 facilitates new investigations of ill-behavior and serves as a challenging benchmark to evaluate downstream tasks.
@@
Justification of Additional Learners.
The Complement Naive Bayes learner was introduced to resolve poor assumptions of the Multinomial Naive Bayes classifier, hence we include it [62]. Complement and Multinomial Naive Bayes are intended for text classification, yet few datasets are text datasets. Therefore, we decided to also include Gaussian Naive Bayes, which assumes features are Gaussian, which can be more reasonable for our diversity of datasets. We, however, choose to include all Naive Bayes learners, as they were also included in the LCDB 1.0. The Nearest Centroid is a classifier with very small training times, and it is known to display ill-behavior in toy settings [29], which is why we included it.
@@ Appendix C Absolute Performance Comparison
Similar to Table 4, we show the statistics of different learners in LCDB 1.1 for both CC-18 and FULL with no data-leakage version in Table 6. Note that non-monotone, non-convex, and ill-behaved refer to shapes that violate monotonicity, convexity, and either, respectively.
Table 6: Statistics of each learner in LCDB 1.1.
@@
SVM_Linear 0.4 3.4 3.8 4.5 5.7 2.3 2.6 0.4 7.9 2.6 2.3 3.4 1.1 1.1 0.0 2.3 3.4 4.2 5.3 1.9 2.3
SVM_Poly 0.4 17.4 3.4 6.4 7.9 3.8 3.4 0.4 6.0 1.5 2.6 3.8 0.8 1.9 0.0 12.8 2.3 5.3 6.0 2.3 1.9
SVM_RBF 0.4 18.5 3.8 15.1 15.8 7.5 2.3 0.4 13.6 3.0 9.4 10.9 4.2 1.9 0.0 12.5 3.4 6.0 7.2 1.5 0.8
SVM_Sigmoid 0.4 19.2 47.9 37.7 58.1 8.3 46.0 0.4 20.0 44.2 34.3 55.5 8.7 45.3 0.0 11.7 39.6 24.9 42.6 0.8 42.6
Decision Tree 0.4 4.5 0.8 1.5 1.5 0.8 1.5 0.4 2.3 1.5 1.5 2.3 0.8 1.5 0.0 3.0 1.5 1.5 2.3 0.8 1.5
ExtraTree 0.4 3.8 1.9 0.4 1.9 0.0 1.1 0.4 2.3 1.9 0.4 1.9 0.0 1.1 0.0 2.6 1.9 0.4 1.9 0.0 1.1
LogisticRegression 0.4 6.8 2.3 4.5 5.3 1.5 1.9 0.4 11.3 1.5 1.5 1.9 1.5 2.3 0.0 7.9 3.4 4.2 4.9 3.0 1.5
PassiveAggressive 0.4 5.7 8.3 4.9 9.4 2.3 3.8 0.4 5.3 2.3 0.8 2.6 0.4 1.9 0.0 1.9 4.2 1.9 4.9 1.1 3.0
Perceptron 0.4 3.0 3.0 2.3 3.8 1.1 1.9 0.4 1.9 1.5 0.8 1.9 0.8 1.1 0.0 1.1 3.0 1.9 3.0 1.1 2.3
RidgeClassifier 0.4 7.2 14.7 15.1 17.0 10.6 5.3 0.4 10.9 2.3 3.8 4.2 1.9 1.9 0.0 6.0 14.7 15.8 17.7 10.2 2.6
SGDClassifier 0.4 2.3 3.0 1.1 3.4 0.8 3.0 0.4 3.0 1.1 1.1 1.5 1.1 2.3 0.0 0.8 3.4 3.0 4.2 1.5 1.9
MLP 0.4 4.9 21.1 21.9 27.9 8.3 4.2 0.4 7.5 18.5 23.0 26.8 6.4 2.3 0.0 3.0 9.4 14.0 16.6 3.0 1.9
LDA 0.4 3.8 32.5 33.2 37.7 24.5 6.8 0.4 4.2 32.5 33.2 37.7 24.5 6.8 0.0 3.0 32.1 33.6 37.7 23.8 7.2
QDA 0.4 3.8 33.6 37.0 45.7 12.1 27.2 0.4 3.4 34.7 40.4 48.3 11.7 30.2 0.0 1.9 38.1 39.6 47.5 9.4 33.6
BernoulliNB 0.4 26.4 11.7 23.0 28.3 13.6 7.9 0.4 21.9 17.4 12.5 23.0 2.6 18.9 0.0 9.1 8.3 8.3 12.1 1.1 9.4
MultinomialNB 30.6 9.1 4.5 6.8 7.9 1.9 6.0 0.4 14.3 14.0 9.8 18.5 2.6 18.9 100.0 0.0 0.0 0.0 0.0 0.0 0.0
ComplementNB 30.6 8.3 6.8 5.7 8.3 1.5 7.5 0.4 5.3 17.7 11.7 19.2 2.6 18.5 100.0 0.0 0.0 0.0 0.0 0.0 0.0
GaussianNB 0.4 4.5 22.6 15.5 24.9 4.2 25.3 0.4 4.5 27.2 19.6 29.1 8.3 26.0 0.0 3.4 28.3 21.1 30.6 8.3 27.5
KNN 0.4 10.9 2.6 1.9 3.8 0.8 4.5 0.4 10.2 2.6 3.0 4.5 0.8 3.0 0.0 10.9 5.7 3.4 6.8 0.8 4.5
NearestCentroid 0.4 10.9 7.2 4.9 7.9 0.8 13.2 0.4 6.8 21.1 9.8 21.5 1.1 27.2 0.0 8.7 12.8 7.5 13.2 1.1 15.5
ens.ExtraTrees 0.4 9.1 1.9 2.3 3.4 1.1 1.9 0.4 8.3 1.9 2.3 3.4 1.1 1.9 0.0 8.3 1.9 2.3 3.4 1.1 1.9
ens.RandomForest 0.4 9.1 1.5 1.9 3.0 1.1 1.5 0.4 8.7 1.5 2.3 3.0 1.1 1.1 0.0 8.3 1.5 2.3 3.0 1.1 1.1
ens.GradientBoosting 0.4 3.4 1.1 0.8 1.9 0.4 0.8 0.4 3.0 1.1 1.1 2.3 0.4 0.8 0.0 3.0 0.8 0.4 1.1 0.4 0.4
DummyClassifier 0.4 69.1 16.6 54.3 60.0 48.7 3.8 0.4 69.1 16.6 54.3 60.0 48.7 3.8 0.0 69.8 16.6 54.7 60.4 49.1 3.8
Appendix E Measure the Shape of Learning Curves
@@ E.1 Surprising Shapes of MLP
E.2 Standard Deviation Distribution of SVM
@@
E.3 More Details regarding LDA
Empirical Observation

### v2

@@ 5.3 DA2: Can Feature Scaling Mitigate Ill-Behavior?
To understand the impact of feature scaling, we now compare the results across scaling techniques. Table 4 indicates that feature scaling marginally reduces the amount of ill-behavior. From Figure 4 we observe that for most learners, feature scaling does not resolve ill-behavior. The Sigmoid SVM becomes slightly more monotone and has fewer peaks, but is still significantly ill-behaved. While preprocessing does not reduce ill-behavior, the SVM absolute performance improves notably and training becomes more stable after scaling is applied (see Appendix E.3 for details). Nearest Centroid and some Naive Bayes models are the only models that become significantly more ill-behaved with feature scaling. Note that GaussianNB is not entirely invariant to feature scaling, due to the way it calculates the variance for numerical stability. The biggest reductions in ill-behavior occur for the Ridge classifier and MLP. Ridge becomes almost completely monotone and without peaks when using min-max scaling, but not with standard scaling. The MLP improves significantly when using standard scaling, largely resolving ill-behaviors, however, min-max scaling does not always help the MLP. We confirm LDA is insensitive to feature scaling, and find the peak occurs when the training set size is approximately equal to the dimensionality (Appendix E.4) in line with peaking literature.
@@ 5.4 DA3: How Do Monotonicity and Convexity Violations Affect Curve Fitting?
In Figure 5 we show how the curve fitting performance is affected by convexity and monotonicity violations. We focus here on the results for the parametric formula $\mathrm{POW4}$ (power law). For MMF4 and WBL4, results are similar, see Appendix F. Performance is measured using the mean squared error (MSE) on the fitted points (interpolation). The mean of the log MSE for monotone curves is over ten times smaller than for non-monotone curves, and the same applies to convex versus non-convex curves. Figure 5 visualizes the MSE versus the violation error. The results reveal a clear positive correlation between violation error and MSE. Our findings clearly show that parametric model fitting is significantly harder for non-monotone and non-convex curves, establishing LCDB 1.1 as a challenging benchmark for evaluating learning curve modeling methods.
@@ 6 Discussion
The shape analysis is challenging. Note that ill-behavior may change if curves are longer (especially dipping). It is also difficult to maintain statistical rigor and consistency; this is because the Bonferroni correction imposes different levels of stringency, for instance, when testing monotonicity (two anchors) and convexity (three anchors). Bonferroni is also quite pessimistic, and some subjective choices had to be made. For example, peaking can also be detected differently (see Appendix H), yet results are similar. Moreover, we have performed additional analysis using slightly less conservative method called Holm’s Step-Down Procedure (Holm’s method) [18], in the sense that it will reject more null hypotheses, typically resulting in fewer Type II errors. This slightly increases the proportion of ill-behaved cases to 19%, but preserves overall consistency (see Appendix I). An analysis using E-values [74] or controlling the false discovery rate [75] may alleviate inconsistency issues, but such an analysis is non-trivial, going beyond our main point.
@@
We have tried our best to make LCDB 1.1 fully reproducible, by using a docker container and fixed python package versions, yet we find that one LDA curve and several QDA curves are non-reproducible, likely due to numerical non-determinism of the singular value decomposition. For this reason, we exclude the QDA learner from the analysis but include its curves in the database.
@@
The next step is to investigate whether ill-behavior persists under hyperparameter tuning, which we leave for future work, since collecting learning curve data with tuning is computationally expensive (LCDB 1.1 already required 800K CPU hours; see Appendix J). This investigation is particularly relevant for models that may exhibit strong sensitivity to hyperparameter settings, such as TabNet. Moreover, hyperparameters in Scikit have reasonable defaults determined by the community, and as such the ill-behavior observed remain surprising and relevant, especially when persistent to different scalings. Although our empirical analysis is scoped to error rate learning curves for classification, the observed ill-behaviors are not confined to this setting. It is therefore valuable to examine alternative evaluation metrics, such as AUC, F1 score, and log-loss, all of which we provide in LCDB 1.1.
7 Conclusion
@@
In conclusion, we introduce the Learning Curves Database 1.1 (LCDB 1.1). This database is more reproducible, of higher resolution, and has multiple types of preprocessing (with and without data-leakage), as well as more modern learners such as CatBoost, TabNet, RealMLP, and TabPFN, making it a valuable database for the community to study learning curves. Moreover, we carefully study ill-behavior and find that a significant amount, 15%, of the curves exhibit ill-behavior while some learners misbehave more frequently than others. Feature scaling rarely solves this problem and in some cases can make it worse. Lastly, we demonstrate the impact of ill-behavior on downstream tasks, underscoring the practical implications. We hope that LCDB 1.1 facilitates new investigations of ill-behavior and serves as a challenging benchmark to evaluate downstream tasks.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge Jesse Krijthe, Taylan Turan, and the anonymous reviewers for their valuable reviews and insightful suggestions. We also thank Jan van Gemert, Marcel Reinders, David Tax, and Gijs van Tulder for their constructive discussions. We further acknowledge the support of DAIC (Delft AI Cluster) and DelftBlue for providing computational resources. Moreover, we thank the community, particularly OpenML for enabling the sharing of data and Scikit-learn for providing essential machine learning tools.
@@
Justification of Other Additional Learners.
The Complement Naive Bayes learner was introduced to resolve poor assumptions of the Multinomial Naive Bayes classifier, hence we include it [76]. Complement and Multinomial Naive Bayes are intended for text classification, yet few datasets are text datasets. Therefore, we decided to also include Gaussian Naive Bayes, which assumes features are Gaussian, which can be more reasonable for our diversity of datasets. We, however, choose to include all Naive Bayes learners, as they were also included in the LCDB 1.0. The Nearest Centroid classifier is computationally efficient but is known to display ill-behavior in toy settings [35].
@@
Modern Learner Implementation Details.
We use the official implementations of CatBoost, TabNet, RealMLP, and TabPFN v2 with their default hyperparameters. For TabNet, we employ the small-scale model (TabNet-S) and use the default hyperparameters without early stopping, different from how TabNet was configured in [23], to ensure consistency across all learners.
@@ Appendix C Absolute Performance Comparison
Furthermore, we can investigate the learners’ optimal performance by filtering datasets according to their characteristics. This analysis can serve as a simple use case that provides further evidence on whether deep learning models outperform tree-based methods under different scale of dataset. Figure 8 compares the learners’ performance between two groups: one where at least one anchor includes more than 10k training samples, and another where all anchors have fewer than 10k training samples. As shown, CatBoost and RealMLP demonstrate consistently strong performance across both groups, while TabNet exhibits competitive performance only on datasets with larger training set sizes.
Refer to caption(a)
Refer to caption(b)
Figure 8: Comparison of learner best performance on average under different feature scaling strategies. (a) LCDB 1.1 FULL (200 datasets with number of samples less than 10k). (b) LCDB 1.1 FULL (65 datasets with number of samples more than 10k).
@@
Similar to Table 4, we show the statistics of different learners in LCDB 1.1 for both CC-18 and FULL with no data-leakage version in Tables 6, 7, and 8. Here, non-monotone, non-convex, and ill-behaved refer to shapes that violate monotonicity, convexity, and either, respectively. In Table 6, we only include the no feature scaling version, since the three considered models can internally handle feature scaling. Note, some datasets from CC-18 or FULL were used in the meta-train benchmark for designing and meta-tuning RealMLP. For Table 7, the relatively high proportion of missing curves is mainly due to the limitations of TabPFN v2, which only supports datasets with up to 10k training samples, 500 features, and 10 classes, as well as curves whose maximum length is below 10k. Moreover, some datasets included in CC-18 and FULL were used during the design of the TabPFN prior. Therefore, the comparison involving TabPFN v2 is not entirely fair, and we present this table separately.
Table 6: Statistics of CatBoost, TabNet, and RealMLP in LCDB 1.1 (no feature scaling).
Learner / Ratio(%) LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
Missing Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Non-Monotone Non-Convex Ill-behaved Peaking Dipping
CatBoost 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.4% 1.1% 1.5% 0.8% 0.4%
TabNet 0.0% 11.1% 72.2% 72.2% 33.3% 1.4% 0.4% 17.4% 73.6% 74.3% 42.3% 4.2%
RealMLP 0.0% 0.0% 1.4% 1.4% 0.0% 0.0% 0.0% 0.4% 4.9% 5.3% 0.0% 0.0%
Table 7: Statistics of TabPFN v2 in LCDB 1.1.
Learner / Ratio(%) LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
no FS min-max FS standardization FS no FS min-max FS standardization FS
Missing 12.5% 12.5% 12.5% 20.8% 20.8% 20.8%
Non-Monotone 0.0% 1.4% 0.0% 0.0% 0.4% 0.4%
Non-Convex 0.0% 1.4% 0.0% 1.5% 2.3% 1.9%
Ill-behaved 0.0% 2.8% 0.0% 1.5% 2.6% 1.9%
Peaking 0.0% 0.0% 0.0% 0.4% 0.4% 0.4%
Dipping 0.0% 0.0% 0.0% 0.0% 0.0% 0.0%
Table 8: Statistics of 24 learners in LCDB 1.1.
@@
SVM_Linear 0.0 3.4 3.8 4.5 5.7 2.3 2.6 0.0 7.9 2.6 2.3 3.4 1.1 1.1 0.0 2.3 3.4 4.2 5.3 1.9 2.3
SVM_Poly 0.0 17.0 3.4 6.4 7.9 3.8 3.4 0.0 5.7 1.5 2.6 3.8 0.8 1.9 0.0 12.8 2.3 5.3 6.0 2.3 1.9
SVM_RBF 0.0 18.1 3.8 15.5 16.2 7.2 2.3 0.0 13.2 3.0 9.8 11.3 4.2 1.9 0.0 12.1 3.4 6.0 7.2 1.5 0.8
SVM_Sigmoid 0.0 19.2 48.3 37.7 58.5 7.9 46.4 0.0 20.0 44.5 34.7 55.8 8.7 45.7 0.0 11.7 39.6 24.9 42.6 0.8 42.6
Decision Tree 0.0 4.5 0.8 1.5 1.5 0.8 1.5 0.0 2.3 1.5 1.5 2.3 0.8 1.5 0.0 3.0 1.5 1.5 2.3 0.8 1.5
ExtraTree 0.0 3.8 1.9 0.4 1.9 0.0 1.1 0.0 2.3 1.9 0.4 1.9 0.0 1.1 0.0 2.6 1.9 0.4 1.9 0.0 1.1
LogisticRegression 0.0 6.8 2.3 4.5 5.3 1.5 1.9 0.0 11.3 1.5 1.5 1.9 1.5 2.3 0.0 7.9 3.4 4.2 4.9 3.0 1.5
PassiveAggressive 0.0 5.7 8.3 4.9 9.4 2.3 3.8 0.0 5.3 2.3 0.8 2.6 0.4 1.9 0.0 1.9 4.2 1.9 4.9 1.1 3.0
Perceptron 0.0 3.0 3.0 2.3 3.8 1.1 1.9 0.0 1.9 1.5 0.8 1.9 0.8 1.1 0.0 1.1 3.0 1.9 3.0 1.1 2.3
RidgeClassifier 0.0 7.2 14.7 15.1 17.0 10.6 5.3 0.0 10.9 2.3 3.8 4.2 1.9 1.9 0.0 6.0 14.7 15.8 17.7 10.2 2.6
SGDClassifier 0.0 2.3 3.0 1.1 3.4 0.8 3.0 0.0 3.0 1.1 1.1 1.5 1.1 2.3 0.0 0.8 3.4 3.0 4.2 1.5 1.9
MLP 0.0 4.9 21.1 21.9 27.9 8.3 4.2 0.0 7.5 18.5 23.0 26.8 6.4 2.3 0.0 3.0 9.4 14.0 16.6 3.0 1.9
LDA 0.0 3.8 32.5 33.2 37.7 24.5 7.2 0.0 4.2 32.5 33.2 37.7 24.5 7.2 0.0 3.0 32.1 33.6 37.7 23.8 7.2
QDA 0.0 3.8 34.0 37.4 46.0 12.1 27.5 0.0 3.4 35.1 40.8 48.7 11.7 30.6 0.0 1.9 38.1 40.0 47.5 9.8 34.0
BernoulliNB 0.0 26.4 12.1 23.4 28.7 13.6 8.3 0.0 22.3 17.7 12.5 23.4 2.6 19.2 0.0 9.1 8.3 8.3 12.1 1.1 9.4
MultinomialNB 30.2 9.1 4.9 7.2 8.3 2.3 6.0 0.0 14.3 14.3 9.8 18.9 2.6 19.6 100.0 0.0 0.0 0.0 0.0 0.0 0.0
ComplementNB 30.2 8.3 6.8 5.7 8.3 1.5 7.5 0.0 5.3 17.7 11.7 19.2 2.6 18.5 100.0 0.0 0.0 0.0 0.0 0.0 0.0
GaussianNB 0.0 4.5 23.0 15.8 25.3 4.5 25.7 0.0 4.5 27.5 20.0 29.4 8.7 26.4 0.0 3.4 28.3 21.1 30.6 8.3 27.5
KNN 0.0 10.9 2.6 1.9 3.8 0.8 4.5 0.0 10.2 2.6 3.0 4.5 0.8 3.0 0.0 10.9 5.7 3.4 6.8 0.8 4.5
NearestCentroid 0.0 10.9 7.5 5.3 8.3 1.1 13.2 0.0 6.8 21.5 10.2 21.9 1.5 27.2 0.0 8.7 12.8 7.5 13.2 1.1 15.5
ens.ExtraTrees 0.0 9.1 1.9 2.3 3.4 1.1 1.9 0.0 8.3 1.9 2.3 3.4 1.1 1.9 0.0 8.3 1.9 2.3 3.4 1.1 1.9
ens.RandomForest 0.0 9.1 1.5 1.9 3.0 1.1 1.5 0.0 8.7 1.5 2.3 3.0 1.1 1.1 0.0 8.3 1.5 2.3 3.0 1.1 1.1
ens.GradientBoosting 0.0 3.4 1.1 0.8 1.9 0.4 0.8 0.0 3.0 1.1 1.1 2.3 0.4 0.8 0.0 3.0 0.8 0.4 1.1 0.4 0.4
DummyClassifier 0.0 69.4 16.6 54.7 60.4 49.1 3.8 0.0 69.4 16.6 54.7 60.4 49.1 3.8 0.0 69.8 16.6 54.7 60.4 49.1 3.8
Appendix E Observation the Shape of Learning Curves
@@ E.1 Surprising Shapes of MLP
E.2 Ill-Behaved Learning Curves in TabNet
The notably ill-behaved learner TabNet exhibits substantial non-convexity in its learning curves, which mostly appears to stem from phase transition phenomena. Figure 12 presents all the ill-behaved learning curves of TabNet in LCDB 1.1 CC-18. We can clearly observe that many of these curves show phase transition shapes, where the performance changes abruptly. From these observations, we find that such ill-behaviors typically occur at small training set sizes. As the training set size increases, the performance of TabNet tends to become more stable. In some cases, multiple phase transitions can even be observed within a single learning curve.
Figure 12: Ill-behaved learning curves of TabNet in LCDB 1.1 CC-18
E.3 Standard Deviation Distribution of SVM
@@
E.4 More Details regarding LDA